OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)? A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (anonymous):

4x - 2y = -6 -4x -4x -2y = -4x - 6 divide by -2 y = 2x + 3 That's parallel, i'm not sure if it goes through (2,-1) though. I haven't checked.